# DLBacktrace: A Model Agnostic Explainability for any Deep Learning Models

### v1

@@ Abstract
The rapid advancement of artificial intelligence has led to increasingly sophisticated deep learning models, which frequently operate as opaque “black boxes” with limited transparency in their decision-making processes. This lack of interpretability presents considerable challenges, especially in high-stakes applications where understanding the rationale behind a model’s outputs is as essential as the outputs themselves. This study addresses the pressing need for interpretability in AI systems, emphasizing its role in fostering trust, ensuring accountability, and promoting responsible deployment in mission-critical fields. To address the interpretability challenge in deep learning, we introduce DLBacktrace, an innovative technique developed by the AryaXAI team to illuminate model decisions across a wide array of domains, including simple Multi Layer Perceptron (MLPs), Convolutional Neural Networks (CNNs), Large Language Models (LLMs), Computer Vision Models, and more.
We provide a comprehensive overview of the DLBacktrace algorithm and present benchmarking results, comparing its performance against established interpretability methods, such as SHAP, LIME, GradCAM, Integrated Gradients, SmoothGrad, and Attention Rollout, using diverse task-based metrics. The proposed DLBacktrace technique is compatible with various model architectures built in PyTorch and TensorFlow, supporting models like Llama 3.2, other NLP architectures such as BERT and LSTMs, computer vision models like ResNet and U-Net, as well as custom deep neural network (DNN) models for tabular data. This flexibility underscores DLBacktrace’s adaptability and effectiveness in enhancing model transparency across a broad spectrum of applications. The library is open-sourced and available at https://github.com/AryaXAI/DLBacktrace.
@@ 1 Introduction
Despite remarkable advancements in artificial intelligence, particularly with the evolution of deep learning architectures, even the most sophisticated models face a persistent challenge: they often operate as "black boxes," with internal processes that remain opaque and difficult to interpret. These models produce highly accurate predictions, yet provide limited insights into how and why they make specific decisions. This opacity raises significant concerns, especially in high-stakes applications like healthcare, finance, and law enforcement, where understanding the rationale behind model outputs is critical. For example, in the healthcare sector, AI-driven diagnostics must be interpretable to ensure that medical professionals can trust and act on recommendations for patient treatment. Similarly, in finance, regulations such as the European Union’s GDPR mandate a "right to explanation" for automated decisions affecting individuals, making explainability not only an ethical imperative but also a regulatory requirement.
The growing demand for explainability and transparency in AI systems is often eclipsed by the prevailing focus on maximizing raw performance. This trend is especially evident with the increasing use of models like OpenAI’s ChatGPT[1], Meta’s LLaMA [2], Google’s Gemini, and similar large language models (LLMs) that are widely adopted for tasks such as language generation, classification, and complex question answering. These models frequently operate as ‘black boxes,’ offering limited insight into their decision-making processes. Many of their architectures are also proprietary and closed-source, adding to their opacity. However, due to their high accuracy and utility in various applications, their lack of transparency is often accepted by users and stakeholders. Consequently, interpretability is frequently overlooked or considered secondary to model performance in deployment. This disregard can lead to challenges, such as increased uncertainty in responses to out-of-domain questions and potential regulatory concerns.
Over the years, various efforts have been made to make machine learning models more transparent, leading to the development of interpretability methods. Popular approaches like Local Interpretable Model-agnostic Explanations (LIME) [3] and SHapley Additive exPlanations (SHAP) [4] aim to clarify model decisions by assigning feature importance scores. These methods work well in tabular data contexts but encounter challenges with complex data types such as images and text, where feature interactions and contextual nuances are intricate. Both LIME and SHAP require repeated evaluations and data perturbations to estimate importance scores, which can increase computational load significantly, especially in high-dimensional contexts. They generate explanations by analyzing selected subsets of data, which leads to instance-specific (local) feature importance. This method may hinder their ability to offer a comprehensive, model-wide perspective, impacting both interpretability and reliability. This issue is particularly pronounced when real-time insights are needed, as the time required to generate explanations increases exponentially with the size of the data subset.
For complex data like images and text, interpretability methods such as Grad-CAM, Integrated Gradients [5], and SmoothGrad help highlight influential regions or tokens in model predictions but come with limitations: Grad-CAM is restricted to CNNs, Integrated Gradients requires carefully chosen baselines, and SmoothGrad’s perturbations add computational cost. For transformer-based models like Vision Transformers (ViTs), techniques such as Attention Rollout aggregate attention across layers to trace information flow, and tools like BertViz [6] visualize attention heads, providing insights into the model’s decision-making. However, interpreting attention alone can be challenging, as attention weights don’t always correlate directly with feature importance, requiring careful analysis for a fuller understanding.
In response to the pressing challenges of interpretability in deep learning, we present, DLBacktrace a model-agnostic method for deep learning interpretability by tracing relevance from output to input, assigning relevance scores across layers to reveal feature importance, information flow, and bias within predictions. Operating independently of auxiliary models or baselines, DLBacktrace ensures deterministic, consistent interpretations across various architectures and data types, including images, text, and tabular data. This approach supports both local (instance-specific) and global (aggregate) analysis, enhancing transparency for models such as LLMs (e.g., BERT, Llama) and computer vision architectures (e.g., ResNet, U-Net), making it a reliable tool for detailed model interpretation and validation. This technique addresses limitations in current interpretability frameworks by delivering stable, reliable explanations that accurately reflect decision-making pathways within a model, with insights at both local (instance-specific) and global (feature-aggregate) levels.
In this work, we make the following contributions:
  • Introduction of DLBacktrace: A detailed methodology outlining the model-agnostic and deterministic approach of DLBacktrace for achieving enhanced interpretability in AI systems.
  • Comprehensive Benchmarking: We benchmark DLBacktrace against widely used interpretability methods (e.g., LIME, SHAP, Grad-CAM, Integrated Gradients and more) across different tasks.
  • Cross-Modality Applications: DLBacktrace’s adaptability is illustrated across various data types, including tabular, image, and text, addressing limitations in current interpretability methods within these domains.
  • Framework for Reliable Interpretability: By providing consistent relevance scores, DLBacktrace contributes to more reliable, regulatory-compliant AI systems, supporting ethical and responsible AI deployment.
@@ 2.1.1 XAI for Responsible and Trustworthy AI
Responsible AI is essential for deploying systems that align with ethical standards and societal values, especially in critical sectors like healthcare, finance, and law enforcement, where AI decisions can profoundly impact individuals and communities. Responsible AI emphasizes fairness, transparency, accountability, privacy, and ethical alignment to prevent bias, protect individual rights, and foster public trust. Fairness ensures that AI models do not discriminate and treat individuals equitably, while transparency provides clear explanations of AI decision-making to build trust. Accountability involves defining responsibility for AI outcomes and enabling human oversight. Privacy and security focus on protecting personal data and ensuring system resilience against security threats, and ethical alignment ensures AI development respects human rights and societal norms.
There has been substantial progress toward responsible AI, supported by regulatory frameworks like the EU’s General Data Protection Regulation (GDPR), which enforces transparency and accountability in automated decision-making. Further advancements include interpretability methods like SHAP, LIME, and Grad-CAM, which aim to make complex models more accessible to stakeholders. Despite this, challenges remain. Current interpretability tools often provide limited insights for complex models, display inconsistencies in high-dimensional data, and are not always applicable in real-time settings. Additionally, the rapid evolution of AI technologies outpaces regulatory responses, creating oversight gaps, especially with advanced models such as large language models (LLMs). These challenges underscore the ongoing need for innovations in responsible AI practices and supportive policy development.
In their work, Madsen et al. [7] advocate for moving beyond traditional interpretability methods, introducing three paradigms that focus on faithful model explanations: models with inherent faithfulness measures, those trained to provide faithful explanations, and self-explaining models. Their work highlights that trustworthy AI requires explanations closely aligned with the model’s actual decision-making to prevent misunderstandings and misplaced trust. Whereas, Singh et al. [8] explore interpretability challenges in large language models (LLMs), proposing tailored methods to address their scale and complexity, enabling clearer insights into model behavior. Tull et al. [9] present a category theory-based framework to unify interpretability approaches, aiming for a cohesive understanding of model behavior crucial to responsible AI. Dinu et al. [10] critically examine assumptions in feature attribution, finding that some methods lack reliability, thus stressing the need for rigorous evaluation of interpretability techniques. While, Kaur et al. [11] apply sensemaking theory to AI, promoting explanations that align with human cognitive processes to enhance trust and accountability by making model reasoning more accessible.
@@ 2.1.2 XAI for Safe AI
AI safety aims to ensure AI systems are predictable, controllable, and aligned with human values, especially in critical areas like healthcare, autonomous vehicles, and infrastructure. A key aspect of AI safety is explainability, as transparent models enable developers, users, and regulators to understand system behavior, detect risks, and implement safeguards. Core pillars of AI safety include robustness, reliability, alignment with human intentions, and the use of explainability to clarify decision processes. Explainability is essential for risk mitigation in dynamic environments, where understanding AI behavior allows for safe intervention. For example, [12] discusses challenges like reward hacking and exploration hazards in reinforcement learning. Explainability techniques help analyze reward structures and behavior patterns, enabling developers to detect and correct unintended actions. Catastrophic forgetting, where models lose prior knowledge when learning new tasks, poses risks in sequential learning [13]. Explainable AI (XAI) methods can identify model areas vulnerable to forgetting, supporting memory mechanisms that preserve safety-critical information. [14] introduces reward modeling to align AI with human preferences through feedback. Explainability tools provide insights into reward structures’ impact on agent behavior, aiding iterative refinement to align model goals with human values. Explainability also supports adversarial robustness by revealing patterns in model vulnerability, guiding defenses that enhance safety and reliability. These examples underscore explainability’s role in AI safety, enhancing transparency, accountability, and risk mitigation. Embedding explainable practices within AI safety frameworks helps developers control AI behavior, reducing risks in diverse operational contexts.
@@ 2.1.3 XAI for Regulatory AI
Explainable AI (XAI) is crucial for regulatory compliance, promoting transparency, fairness, and accountability in AI-driven decisions in sectors like finance, healthcare, and law. Regulatory frameworks increasingly mandate interpretable models to ensure oversight, protect user rights, and uphold ethical standards. Key elements of XAI in regulatory contexts include transparent decision processes, model auditability, and mechanisms to mitigate bias. In finance, XAI helps institutions clarify decisions on credit scoring, loan approvals, and fraud detection, reducing regulatory risks and fostering public trust. For instance, [15] systematically reviews XAI applications in finance, illustrating transparency’s role in regulatory compliance. As large language models (LLMs) become ubiquitous, interpretability in NLP has gained importance. [16] examines alignment of interpretability methods with stakeholder needs, categorizing techniques and identifying differences between developer and non-developer requirements. Stakeholder-centered frameworks help ensure more responsible AI deployment. In healthcare, XAI is vital for patient safety and ethical standards. Explainable models enable healthcare providers to interpret AI-driven diagnoses, treatments, and risk assessments, aligning these with medical regulations. [17] addresses the challenge of conflicting post hoc explanations, often resolved by practitioners ad hoc. This study calls for standardized metrics to enhance explanation reliability, especially in high-stakes fields like healthcare and finance. Specific to LLMs, [8] explores their potential for interactive, natural language explanations that can improve comprehension of complex behaviors. The authors address interpretability challenges, such as hallucinations and computational cost, recommending LLM-based methods to improve transparency and nuanced insights in high-accountability domains. Finally, [18] critiques common interpretability techniques, highlighting limitations in methods like Layer-wise Relevance Propagation (LRP) [19] and proposing Cosine Similarity Convergence (CSC) as a metric to improve explanation accuracy. These studies emphasize XAI’s role in enhancing regulatory compliance by fostering transparency, accountability, and fairness. Embedding explainable practices in regulatory frameworks ensures ethical AI use, facilitates oversight, and builds trust across regulated sectors.
@@ 2.2.1 Tabular Data
In high-stakes applications such as finance and healthcare, machine learning models like regression and probabilistic algorithms (e.g., decision trees and their variants) are often preferred. This is because deep learning models are frequently described as "black boxes," making it challenging to interpret how they arrive at their conclusions. This lack of transparency can be particularly problematic in critical contexts where accountability and trust are essential.
To enhance interpretability, explainable algorithms like LIME [3] and SHAP [4] are increasingly used. LIME (Local Interpretable Model-Agnostic Explanations) builds simple, interpretable models around specific data points to highlight the most influential features for each prediction. SHAP (SHapley Additive exPlanations) assigns importance scores to each feature, providing both global explanations (by ranking features based on overall importance) and local explanations (by illustrating how individual features contribute to specific predictions).
However, these methods have limitations. LIME’s reliance on random sampling can lead to inconsistent explanations for the same data point, and its effectiveness can be sensitive to the choice of perturbation method. SHAP, while comprehensive, can be computationally expensive for large datasets and complex models. Additionally, SHAP’s model-agnostic nature may result in less accurate explanations for highly intricate models, like deep neural networks. As a result, both LIME and SHAP may face challenges in providing precise, interpretable explanations for complex deep learning models.
@@ 2.2.2 Image Data
In explainable AI (XAI) for image modality-based tasks, gradient-based methods such as GradCAM [20], Vanilla Gradient [21], SmoothGrad [22], and Integrated Gradients [5] are widely used for interpreting model predictions. GradCAM generates heatmaps by calculating the gradient of the target class with respect to convolutional layer activations, but may miss fine details and is sensitive to input noise. Vanilla Gradient directly computes gradients on the input, though it faces the "saturation problem," where gradients become too small for clear interpretation. SmoothGrad improves clarity by averaging gradients with added noise, albeit at a computational cost, while Integrated Gradients addresses saturation by calculating an integral of gradients from a baseline to the input, though it also demands significant computation.
In recent advances, Vision Transformers (ViTs) require specific interpretability approaches due to their reliance on attention mechanisms. In the paper [23], authors introduces TokenTM, a method designed for ViTs that considers both token transformations (changes in token length and direction) and attention weights across layers. By aggregating these factors, TokenTM provides more focused and reliable explanations, addressing unique interpretability challenges in transformer models. These developments reflect a shift in XAI, where interpretability techniques are tailored to the unique demands of model architectures, like CNNs and transformers, enhancing transparency and reliability across different models.
@@ 2.2.3 Textual Data
In the text modality, quite a few explanation methods are employed to enhance the interpretability of machine learning models. Among these, LIME [3] and SHAP [4] are baseline for interpreting text classification models. Gradient-based methods, such as GradCAM [20], Integrated Gradients [5], and Attention Rollout [24], also play a significant role across diverse model architectures.
For text generation tasks, 17 challenges were identified by [25], such as tokenization effects and randomness, and advocates for probabilistic explanations and perturbed benchmarks to address these issues. Furthermore, LACOAT introduced by [26], which clusters word representations to produce context-aware explanations. It maps test features to latent clusters and translates these into natural language summaries, improving interpretability for complex NLP tasks. While [27] draws attention to the noise in explanations generated by large language models (LLMs), noting significant randomness. The study suggests that more sophisticated interpretability techniques, beyond simple word-level explanations, are needed to achieve reliability. Lastly, [28] proposes a hybrid approach that combines counterfactual explanations with domain knowledge. This method generates context-specific counterfactuals and incorporates user feedback, enhancing BERT’s interpretability through expert insights and interactive elements, thus fostering a more transparent framework.
In addition to the previously discussed methods, Mechanistic Interpretability has become a cornerstone of research aimed at unraveling the internal mechanisms of large language models (LLMs). This field focuses on dissecting how specific components, such as neurons and attention heads, contribute to a model’s functionality and decision-making processes. For example, Olah et al. [29] performed an in-depth analysis of individual neurons in LLMs, revealing how certain neurons specialize in detecting specific linguistic features. Their work highlights the modularity and specialization inherent in these models. Building on this foundation, Elhage et al. [30] introduced the concept of "induction heads" in transformer architectures. They demonstrated that these attention heads play a critical role in tasks such as sequence copying, offering insight into the mechanisms underlying certain model behaviors. Further advancing the field, Nanda et al. [31] investigated the phenomenon of "grokking" in LLMs, where models suddenly exhibit strong generalization capabilities after extended training. Their analysis traced the internal changes leading to this abrupt performance improvement, shedding light on this intriguing aspect of model behavior. These advancements underscore the significance of mechanistic interpretability in enhancing our understanding of the intricate operations of LLMs. By making AI systems more transparent and reliable, this research is vital for fostering trust, particularly as LLMs are scaled and deployed in high-stakes applications.
@@ 2.2.4 Metrics for Benchmarking Explainability
Recent advancements in Explainable AI (XAI) emphasize the need for robust evaluation frameworks, benchmarks, and specialized toolkits to enhance transparency and trust in machine learning systems. Quantus [32], provides a modular toolkit with metrics such as faithfulness, robustness, and completeness, promoting reproducible and responsible evaluation of neural network explanations. BEExAI [33] addresses the underexplored domain of tabular data by benchmarking explanation methods using tailored datasets and metrics like local fidelity and human interpretability, enabling systematic comparisons in real-world scenarios. In a survey over 30 XAI toolkits by [34], including Quantus, highlighting challenges such as inconsistent metrics, lack of fairness assessments, and insufficient focus on human-centered evaluation, calling for standardized benchmarks and unified platforms. PyXAI [35] complements these efforts by focusing on tree-based models, introducing efficient tools and unique metrics such as path importance and rule-level interpretability for domains like healthcare and finance. Despite these advancements, key gaps remain, including standardization across tools, evaluation for diverse data types, human-centered usability, and fairness and robustness assessments, underscoring the need for continued research to achieve actionable and responsible XAI.
3 Backtrace

### v2

@@ Abstract
The rapid growth of AI has led to more complex deep learning models, often operating as opaque "black boxes" with limited transparency in their decision-making. This lack of interpretability poses challenges, especially in high-stakes applications where understanding model output is crucial. This work highlights the importance of interpretability in fostering trust, accountability, and responsible deployment. To address these challenges, we introduce DLBacktrace, a novel, model-agnostic technique designed to provide clear insights into deep learning model decisions across a wide range of domains and architectures, including MLPs, CNNs, and Transformer-based LLM models. We present a comprehensive overview of DLBacktrace and benchmark its performance against established interpretability methods such as SHAP, LIME, and GradCAM. Our results demonstrate that DLBacktrace effectively enhances understanding of model behavior across diverse tasks. DLBacktrace is compatible with models developed in both PyTorch and TensorFlow, supporting architectures such as BERT, ResNet, U-Net, and custom DNNs for tabular data. The library is open-sourced and available at https://github.com/AryaXAI/DLBacktrace.
@@ 1 Introduction
Deep learning models, including large language models (LLM) such as ChatGPT [1], LLaMA [2], and Google’s Gemini, have achieved remarkable performance in various applications such as language generation, image recognition, and complex reasoning tasks. However, these sophisticated systems remain largely opaque, offering limited insight into the decision-making processes that underpin their predictions. This opacity poses significant challenges, particularly in high-stakes domains such as healthcare, finance, and law enforcement. In these areas, understanding model rationales is not merely desirable—it is often a regulatory and ethical imperative. For example, healthcare professionals must trust and understand AI-assisted diagnostics before implementing treatments, and regulatory frameworks such as the EU GDPR (Article 22) mandate explainability in automated decisions that affect individuals.
The demand for responsible and trustworthy AI has led to the rise of eXplainable AI (XAI) research, aiming to align models with ethical standards, ensure fairness, transparency, and accountability, and increase public trust [3, 4, 5, 6]. Similarly, AI safety research underscores that explainability helps identify and mitigate risks such as reward hacking and catastrophic forgetting, guiding the development of robust systems [7, 8, 9]. Various methods have emerged to address interpretability. Model-agnostic approaches like LIME [10] and SHAP [11] assign feature importance scores. However, these are often costly, sensitive to data perturbations, and struggle with high-dimensional data such as images and text. Similarly, gradient-based methods like Grad-CAM and Integrated Gradients [12], can highlight influential regions or tokens, but face architectural constraints, dependence on carefully chosen baselines, and complex interpretation of attention weights.
Interpretability frameworks often lack unified metrics, complicating comparisons, and sometimes provide conflicting explanations, underscoring the need for standardized evaluation criteria [13, 14, 15]. Compliance with regulatory standards amplifies the importance of transparency. In finance, explainable credit scoring, loan approvals, and fraud detection are increasingly mandated [16]. In healthcare, explainable models ensure the trust needed to integrate AI-driven diagnostics responsibly. The complexity of LLMs, in particular, necessitates tailored interpretability strategies to address challenges like hallucinations, computational constraints, and scale, bringing user-centered methods to the forefront [17].
To address the growing need for effective interpretability methods in deep learning, we present DLBacktrace, a model-agnostic approach that works independently of auxiliary models or baselines. DLBacktrace ensures consistent and reliable explanations across various architectures and data types, including images, text, and tabular data. By tracing relevance from the output back to the input, it assigns relevance scores across layers to highlight feature importance, information flow, and potential biases in predictions. In this work, we make the following contributions:
  • We introduce DLBacktrace, a robust, model-agnostic method for deep learning interpretability.
  • We provide a detailed explanation of the approach behind DLBacktrace, addressing challenges related to stability and consistency across different model architectures.
  • We demonstrate the applicability of DLBacktrace across different domains, including text, image, and tabular data, paving the way for its broader adoption in responsible AI applications.
Figure 1: DLBacktrace Workflow: Generating Fine-Grained Explanations from Pre-Trained Models and Test Time Instance.
@@ 2.1.1 XAI for Responsible and Trustworthy AI
Responsible AI ensures AI systems align with ethical standards, particularly in sectors like healthcare, finance, and law enforcement, where decisions impact individuals and communities. It focuses on fairness, transparency, accountability, privacy, and ethical alignment to minimize bias, protect rights, and build trust. While progress has been made through regulations like the EU’s GDPR and tools such as SHAP, LIME, and Grad-CAM, challenges remain. Current methods provide limited insights, especially with high-dimensional data, and are not always suitable for real-time use. Moreover, rapid AI advancements often outpace regulatory efforts, creating oversight gaps, particularly with large language models (LLMs). [5] advocate for moving beyond traditional interpretability techniques, proposing three approaches that focus on faithful model explanations: models with inherent faithfulness measures, those trained to offer faithful explanations, and self-explaining models. [3] address interpretability challenges in LLMs, offering methods tailored to their scale and complexity. [6] introduce a category theory-based framework to unify interpretability approaches. [14] critique feature attribution methods, highlighting the need for more reliable evaluations. [4] apply sensemaking theory to AI, advocating for explanations that align with human cognitive processes to improve trust and accountability.
@@ 2.1.2 XAI for Safe AI
AI safety ensures that AI systems are predictable, controllable, and aligned with human values, particularly in high-stakes areas like healthcare, autonomous vehicles, and infrastructure. A key aspect of AI safety is explainability, which allows developers, users, and regulators to understand system behavior, detect risks, and implement safeguards. Core principles of AI safety include robustness, reliability, alignment with human intentions, and the use of explainability to clarify decision-making. Explainability is vital for risk mitigation in dynamic environments, where understanding AI behavior enables safe interventions. For instance, [7] highlights challenges like reward hacking and exploration hazards in reinforcement learning, where explainability techniques help analyze reward structures and behavior patterns to prevent unintended actions. Catastrophic forgetting, where models forget prior knowledge when learning new tasks, presents risks in sequential learning [8]. XAI methods can identify vulnerable areas in models, supporting mechanisms that preserve critical information. Explainability also aids in improving adversarial robustness by revealing vulnerabilities, guiding defenses that enhance safety.
@@ 2.1.3 XAI for Regulatory AI
Explainable AI (XAI) is essential for regulatory compliance, promoting transparency, fairness, and accountability in AI-driven decisions across sectors like finance, healthcare, and law. Increasingly, regulatory frameworks require interpretable models to ensure oversight, protect user rights, and uphold ethical standards. Key components of XAI in regulatory contexts include transparent decision-making, model auditability, and bias mitigation. In finance, XAI helps clarify decisions on credit scoring, loan approvals, and fraud detection, reducing regulatory risks and building public trust. [16] reviews XAI applications in finance, emphasizing transparency’s role in regulatory compliance. As large language models (LLMs) gain prominence, interpretability in NLP becomes crucial. [17] examines how interpretability methods align with stakeholder needs, highlighting the importance of stakeholder-centered frameworks for responsible AI deployment. In healthcare, XAI supports patient safety and ethical standards by helping providers interpret AI-driven diagnoses and treatments. [3] explores how LLMs can provide interactive, natural language explanations to improve understanding of complex AI behaviors, addressing challenges like hallucinations and computational costs. These studies underscore XAI’s role in enhancing regulatory compliance by fostering transparency, accountability, and fairness, ensuring ethical AI use and facilitating oversight across regulated sectors.
@@ 2.2.1 Tabular Data
In high-stakes fields like finance and healthcare, machine learning models such as regression and decision trees are often preferred over deep learning models, which are seen as "black boxes." Explainable algorithms like LIME [10] and SHAP [11] address this by enhancing interpretability. LIME creates simple models around specific data points to highlight key features, while SHAP assigns importance scores to features, providing both global and local explanations. However, both methods have limitations. LIME’s random sampling can lead to inconsistent explanations, and SHAP can be computationally expensive for large datasets. Additionally, SHAP’s model-agnostic approach may be less effective for complex Deep Neural Networks.
@@ 2.2.2 Image Data
For image data, gradient-based methods such as GradCAM [18], Vanilla Gradient [19], SmoothGrad [20], and Integrated Gradients [12] are widely used to interpret model predictions. GradCAM generates heatmaps based on gradients, while Vanilla Gradient faces the "saturation problem." SmoothGrad improves clarity by adding noise to gradients, and Integrated Gradients address saturation but require more computation.
@@ 2.2.3 Textual Data
For text-based tasks, methods like LIME [10] and SHAP [11] are common for interpreting text classification models, alongside gradient-based methods like GradCAM [18] and Integrated Gradients [12]. For text generation, challenges such as tokenization effects and randomness are noted by [21], and techniques like LACOAT [22] improve interpretability by clustering word representations. [23] introduces a soft-erasure method to improve faithfulness in NLP explanations, while [24] proposes Normalized AOPC to standardize feature attribution metrics. Mechanistic interpretability is gaining traction in LLMs, focusing on understanding how individual components, such as neurons and attention heads, contribute to decision-making [25, 26].
@@ 2.2.4 Metrics for Benchmarking Explainability
Recent work has focused on developing robust frameworks for evaluating the quality of XAI methods. Quantus [27] provides a modular toolkit for evaluating neural network explanations, while BEExAI [28] focuses on tabular data, offering metrics like local fidelity. Despite these advancements, challenges like standardization, usability, and fairness remain, highlighting the need for continued research in XAI evaluation.
(a) Sample Network
(b) Relevance for Sample Network
Figure 2: Illustration Depicting DLBacktrace Calculation for a Sample Network
3 DLBacktrace